# Observational Probes of the Neutron Star Equation of State with Hyperons, Bosonic Dark Matter, and Quark Matter

### v2

@@
11institutetext: Institute for Theoretical Physics, University of Wrocław, Plac Maksa Borna 9, 50-204 Wrocław, Poland 22institutetext: Incubator of Scientific Excellence—Centre for Simulations of Superdense Fluids, University of Wrocław, Plac Maksa Borna 9, 50-204 Wrocław, Poland 33institutetext: Frankfurt Institute for Advanced Studies, Giersch Science Center, D-60438 Frankfurt am Main, Germany 44institutetext: Alikhanyan National Science Laboratory (Yerevan Institute of Physics), Alikhanyan Brothers St 2, 0036 Yerevan, Armenia 55institutetext: Technische Universität Darmstadt, Fachbereich Physik, Institut für Kernphysik, Schlossgartenstraße 9, D-64289 Darmstadt, Germany 66institutetext: GSI Helmholtzzentrum für Schwerionenforschung GmbH, Theorie, Planckstraße 1, D-64291 Darmstadt, Germany
@@
Mahboubeh Shahrbaf email: mahboubeh.shahrbafmotlagh@uwr.edu.pl112233    Davood Rafiei Karkevandi email: Davood.Rafiei64@gmail.com11    Alexander Ayriyan email: alexander.ayriyan@uwr.edu.pl1144    Stefan Typel email: stypel@ikp.tu-darmstadt.de5566
@@
The presence of dark matter in neutron stars is of growing interest due to its possible impact on their structure and observables. Among proposed candidates, the hypothetical sexaquark has emerged as a promising bosonic dark matter particle, potentially forming under extreme conditions in neutron star cores. We investigate this scenario using a relativistic density functional approach, including hyperons (DD2Y-T model) for the hadronic phase, coupled to a deconfined quark phase described by a non-local Nambu-Jona-Lasinio model. The phase transition is modeled through a smooth crossover, and sexaquark-baryon interactions are introduced via an effective mass shift representing repulsion. We assess whether this hybrid scenario, incorporating hyperons, bosonic dark matter, and deconfined quark matter, is consistent with multi-messenger constraints from neutron stars. By scanning the parameter space, we identify sexaquark masses that are consistent with mass-radius measurements and tidal deformability. Our results show that the presence of the sexaquark softens the equation of state, enabling the model to satisfy both the radius and tidal deformability constraints around the canonical $1.4M_{\odot}$ neutron stars. We incorporate all available NICER data, including PSR J0437-4715 and newly published PSR J0614-3329. We find out that hybrid stars with sexaquark mass near 1900 MeV agree with all current observational limits, including HESS J1731-347 and PSR J0952-0607, representing the lightest and most massive known neutron stars. Furthermore, we perform a Bayesian analysis, which yields a favored sexaquark mass range of 1885-1935 MeV, supporting the potential role of this exotic particle in the interiors of neutron stars.
@@
Dark matter – Neutron star – Equation of State – Phase transition – Sexaquark – Bayesian Analysis
@@ 1 Introduction
The accepted cosmological models that are presented and founded in General Relativity (GR), suggest that almost 27% of the matter-energy composition of the universe exists in the form of enigmatic Dark Matter (DM) (Bertone & Hooper, 2018; de Martino et al., 2020).
@@
Multi-messenger astrophysics is actively engaged in the search for and examination of new ideas concerning DM particles. These candidates are believed to be non-relativistic particles that exist beyond the Standard Model of particle physics. They interact weakly with ordinary baryonic matter (BM) solely through gravitational forces (Salucci et al., 2021; Arbey & Mahmoudi, 2021). In this context, weakly interacting massive particles (WIMPs) are one of the notable DM candidates (Bertone & Tait, 2018; Arcadi et al., 2018). Moreover, hypothetical bosonic particles such as the axion (Marsh, 2016; Chadha-Day et al., 2022) and the sexaquark (S) (Farrar, 2022; Doser et al., 2023; Moore & Slatyer, 2024) also exhibit the primary characteristics required to be considered as potential DM candidates. Their interaction with the BM must be very weak, which is why they have remained undetected thus far. A broad spectrum for the mass of candidates for DM particles exists, spanning a wide range from approximately $10^{-22}$ eV to $10^{10}$ eV, based on the particular type of candidate (Cirelli et al., 2024). Moreover, their self-interaction is another key parameter of the DM model. Indeed, self-interacting DM has the potential to address several issues in cosmology and astrophysics, including ”core-cusp”, ”missing satellite”, and ”too big to fail” problems (Spergel & Steinhardt, 2000; Peter et al., 2013; Kaplinghat et al., 2016). Given these model parameters, DM particles must be incorporated into an EOS, which is a fundamental tool for characterizing matter.
@@
It is believed that gravitationally stable astrophysical objects composed of DM could form in the universe (Liebling & Palenzuela, 2023; Baryakhtar et al., 2022; Visinelli, 2021). In this context, they can exist as an object made entirely of DM so-called dark stars (Narain et al., 2006; Maselli et al., 2017) such as axion stars, boson stars (Eby et al., 2016; Visinelli et al., 2018; Pitz & Schaffner-Bielich, 2023), or one which has a fraction of DM mixed with BM (Bramante & Raj, 2024). For the latter scenario, compact objects such as black holes (Shakeri & Hajkarim, 2023), neutron stars (NSs) (Karkevandi et al., 2022; Shakeri & Karkevandi, 2024), white dwarfs (Ryan & Radice, 2022) and even hypothetical quark (strange) stars (Ferreira & Fraga, 2023; Yang & Pi, 2024; Sedaghat et al., 2025; Liu et al., 2025a) offer a significant opportunity to probe DM indirectly. These objects are considered promising astrophysical environments where DM may reside, either within them or in a surrounding halo.
@@
NSs, due to their extreme gravitational potential and high baryon density, have the potential to accumulate DM particles and are widely regarded as astrophysical laboratories for studying DM (Grippa et al., 2025). In recent years, motivated by new and significant observational data on NS properties, numerous studies have investigated the possible presence of DM within these objects (Ivanytskyi et al., 2020; Rafiei Karkevandi et al., 2021; Dengler et al., 2022; Rutherford et al., 2023; Karkevandi et al., 2024; Thakur et al., 2024a; Konstantinou, 2024; Routaray et al., 2025; Biesdorf et al., 2025; Arvikar et al., 2025). The influence of DM on observable features of NSs has been explored to place constraints on the parameter space of DM models (Sedrakian, 2016, 2019; Shakeri & Karkevandi, 2024; Giangrandi et al., 2023; Diedrichs et al., 2023; Ávila et al., 2024; Issifu et al., 2025; Cipriani et al., 2025; Scordino & Bombaci, 2025; Mukherjee et al., 2025). Furthermore, unusual measurements of NS mass and radius have prompted attempts to explain them in terms of mixed configurations, hereafter referred to as DM admixed NSs. In this direction, HESS J1731-347 (Sagun et al., 2023; Marzola et al., 2025) and XTE J1814-338 ultra-compact stars (Pitz & Schaffner-Bielich, 2025; Yang et al., 2025; Lopes & Issifu, 2025), as well as very massive compact objects such as the secondary ones detected in GW190814 and GW230529 events, are noteworthy cases (Barbat et al., 2024; Vikiaris et al., 2024).
@@
Concerning the formation of DM admixed NSs, several scenarios are considered, each leading to various DM fractions within the star (Karkevandi et al., 2022). Among these, DM production inside NSs or during supernova explosions is notable mechanism (Ellis et al., 2018; Nelson et al., 2019). An additional neutron decay channel into DM, potentially addressing the neutron decay anomaly, could result in the production of a significant fraction of fermionic DM inside NSs (Shirke et al., 2023, 2024; Thakur et al., 2024b; Das & Burgio, 2025). Furthermore, as a novel concept, the production of S, a deeply bound bosonic DM candidate, offers a promising mechanism for forming these mixed objects (Shahrbaf et al., 2022; Blaschke et al., 2022; Shahrbaf, 2023), which is the focus of this paper.
@@
Apart from the formation of these objects, modeling and investigating DM admixed NSs are done based on the fact that DM and BM interact only through gravitational force, or there is a non-gravitational interaction between them. For the former, two-fluid Tolman-Oppenheimer-Volkof (TOV) equations describe the mixed object (Rafiei Karkevandi et al., 2021; Collier et al., 2022; Karkevandi et al., 2024; Rutherford et al., 2024; Shawqi & Morsink, 2024; Dengler et al., 2025; Liu et al., 2025b). Depending on the DM parameters, it can reside as a core inside NS or may form an extended halo around it. For this approach, one can refer to the works done by Karkevandi et al. (2022); Shakeri & Karkevandi (2024) and the references therein. However, when a non-gravitational interaction is considered between BM and DM, single fluid TOV equations are employed to describe the hydrostatic equilibrium of the object (Lenzi et al., 2023; Shahrbaf, 2023; Guha & Sen, 2024; Hajkarim et al., 2024; Kumar & Sotani, 2025; Klangburam & Pongkitivanichkul, 2025; Das et al., 2025). Therefore, DM production inside NSs, such as S, is investigated using the single-fluid approach, due to the non-gravitational interaction between DM particles and the baryonic medium (Shahrbaf et al., 2022). In this paper, we demonstrate how the presence of S affects the EOS, leading to changes in the observable properties of DM admixed NSs.
@@
It is worth mentioning that the EOS for strongly interacting matter in NSs is tightly bound by observational constraints. Therefore, efficient boundaries on the features of DM candidates have been established using recent advancements in observational astronomy, mainly facilitated by gravitational wave (GW) interferometers (Abbott et al., 2018; Dietrich et al., 2020; Pang et al., 2023; Ecker et al., 2025) and the Neutron Star Interior Composition ExploreR (NICER) (Miller et al., 2019, 2021; Riley et al., 2019, 2021; Vinciguerra et al., 2024; Salmi et al., 2024b, a; Dittmann et al., 2024a; Choudhury et al., 2024b; Mauviard et al., 2025b). In this study, mass, radius, and tidal deformability serve as the fundamental constraints used to investigate these DM admixed objects. Thus, we focus on high-precision mass–radius (M-R) measurements obtained through X-ray pulse profile modeling by the NICER telescope. We include the recent result for PSR J0437–4715 which reports a mass of $M=1.418\pm 0.037M_{\odot}$ and a radius of $R=11.36^{+0.95}_{-0.63}$ km Choudhury et al. (2024b). Importantly, we also incorporate the latest NICER measurement of PSR J0614-3329, which infers a radius of $R=10.29^{+1.01}_{-0.86}$ km for a mass of $M=1.44^{+0.06}_{-0.07}M_{\odot}$ (Mauviard et al., 2025b). These new NICER results are important, as they support a softer EOS and add valuable boundary conditions for constraining models of dense matter in NSs.
@@
Moreover, the detection of GW signals arising from binary NS mergers, along with the subsequent determination of the tidal deformability parameter, have opened a new window for investigating the exotic internal structures of NSs (Abbott et al., 2017, 2020). Therefore, the tidal deformability constraint from the GW170817 event (Abbott et al., 2018) $\Lambda_{1.4}=190^{+390}_{-120}$, is also considered to obtain the allowed parameter space of the S bosonic DM model.
@@
The concept of a deeply bound S, an alternative term for the hexaquark family of hypothetical particles, was first introduced as a bosonic candidate for DM in references Farrar (2003, 2018, 2022); Doser et al. (2023). The quark composition of S is identical to that of the H-dibaryon (Jaffe, 1977), with both consisting of uuddss quarks. However, S exhibits significantly stronger binding compared to the H-dibaryon. If S represents a molecular state like $\Lambda\Lambda$, the binding of two $\Lambda$ particles can exclusively occur through the exchange of color-neutral entities, such as mesons. However, when S comprises a complex system consisting of three colored diquarks, these entities engage in interactions primarily governed by the color force. This force is notably stronger than the meson exchange force, particularly when considering short-range distances.
@@
The proposed S is a boson at a spin-color-flavor singlet and even parity state, characterized by a zero electric charge, a baryon number of 2, and a strangeness value of -2. The potential of the S particle to serve as DM depends on several parameters, such as its mass and its stability against dissociation into two baryons, which requires a small dissociation amplitude. Other important factors include its elastic scattering cross section with nucleons and its annihilation with antibaryons (or antisexaquarks) into baryons or other Standard Model final states. S may manifest itself as a deeply bound state, displaying a mass sufficiently low to be either absolutely or effectively stable against decay. Our main focus in this study is the mass of the S. Notably, all measurements show that its mass is less than 2230 MeV, the threshold for producing two $\Lambda$ baryons. This suggests that the S may be a bound state (Jaffe, 1977; Kodama et al., 1994; Azizi et al., 2020; Gross et al., 2018; Shahrbaf et al., 2022; Evans & Ward, 2023).
@@
However, Lattic QCD calculations suggest that H-dibaryon is an unbound particle (Inoue et al., 2011; Beane et al., 2011; Green et al., 2021; Shanahan et al., 2011), but Farrar (2022) assumes S as a deeply-bound state which is difficult to be detected in lattice calculations, setting it different from the loosely bound H-dibaryon. For the S to be a viable DM candidate, both its possible decay channels and any Standard Model particle decays involving the S must be either kinematically forbidden or highly suppressed. If the S particle were to decay, the energetically most favorable channel would produce two protons and two electrons, with a total mass of approximately $1878$ MeV. Consequently, the S particle would be absolutely stable if its mass lies below this threshold, i.e., $m_{\text{S}}<1878$ MeV. From an experimental point of view, the existence of S with a mass between $2054$ MeV and $2223.7$ MeV has been ruled out (Takahashi et al., 2001). Additionally, Farrar & Wang (2023) argues that if the S mass equals the combined mass of a lambda hyperon, a proton, and an electron ($m_{\text{S}}=2054$ MeV), its decay into two nucleons would involve a doubly weak interaction, resulting in a lifetime longer than the age of the universe. The constraints on the mass of the S have been extensively discussed in Farrar & Wang (2023) and the references therein. Meanwhile, Gal (2024) demonstrated that the observation of double-$\Lambda$ hypernuclei decaying into single-$\Lambda$ hypernuclei does not exclude the existence of a deeply bound S within the mass range of $2m_{n}$ to $m_{\Lambda}+m_{n}$. However, its $\Delta S=2$ decay lifetime ($\tau(H\rightarrow nn)$) is estimated to be shorter than the timescale required for it to be a viable DM candidate. Moreover, the ratio of densities between S DM and ordinary BM in the Universe, which can be estimated through considerations of statistical equilibrium within the quark-gluon plasma, aligns with observational data (Farrar, 2022). Under the above-mentioned conditions, we assume that the S has the potential to be a DM candidate and examine its mass based on observational constraints from NSs. A mechanism for producing such a uuddss state at rest through the formation of antiprotonic atoms has been proposed in Doser et al. (2023). Ultimately, lattice QCD can be used to study the interactions and decay channels of the S, offering deeper insight into its properties.
@@
Complementary to the S DM hypothesis, heavy-ion collision experiments have demonstrated that hyperons can be abundantly produced in high-density environments. These findings are significant for understanding the behavior of dense nuclear matter, as they provide empirical evidence supporting the inclusion of hyperons in the EOS of matter under extreme conditions (Abelev et al., 2010; Rappold et al., 2015; Acharya et al., 2022; Leung, 2022; Acharya et al., 2023). Therefore, including hyperons in NS EOS models is essential for a comprehensive understanding of their structure, stability, and evolution, aligning theoretical predictions with observational data. Moreover, given that our DM candidate carries double strangeness, it is natural to include hyperons in the hadronic phase as well. In this study, the S production is established using a generalized relativistic density functional (GRDF) approach known as DD2Y-T Stone et al. (2021). This hadronic model contains all $J=1/2$ octet baryons, namely the neutron, proton, $\Lambda$, $\Sigma$, $\Xi$ hyperons, giving a more reliable description of the high-density matter inside NSs. The EOS for the entire NS, including both the core and the crust, is constructed consistently using the DD2 crust model as in Typel & Shlomo (2024), which ensures a smooth and thermodynamically consistent matching between the outer layers and the dense core. In addition to the strongly interacting baryonic components, the EOS must also account for leptonic degrees of freedom - specifically, electrons and muons - due to the requirements of charge neutrality and beta equilibrium in NS matter. These leptons are treated as non-interacting, relativistic Fermi gases and play a crucial role in balancing the electric charge introduced by the presence of protons and charged hyperons.
@@
Furthermore, from a phenomenological point of view, a peak in the square of the speed of sound in QCD matter has been observed, which is associated with the percolation threshold and the phase transition to deconfined QM (Satz, 1998; Magas & Satz, 2003; Castorina et al., 2009; Braun-Munzinger et al., 2015; Fukushima et al., 2020; Marczenko et al., 2023). These findings support the idea that the core of NSs may consist not only of hadronic matter but also deconfined QM. As a result, it is highly possible to encounter a phase transition to QM in the core of NS (Li et al., 2023; Gholami et al., 2025; Li et al., 2025; Lindblom et al., 2025). Therefore, our model is connected to an advanced QM model named non local Nambu-Jona-Lasinio (nl-NJL) model based on the crossover phase transition by a novel method.
Constructing a phase transition to deconfined QM and assuming a positive mass shift for S, ensures the $2M_{\odot}$ mass limit for NSs. Notably, the onset of S softens the EOS and thus, it contributes to the consistency of the tidal deformability and radius of NSs around $1.4M_{\odot}$ with observational constraints. As a result, our model provides a hybrid star composed of a dense core of deconfined QM surrounded by layers of DM, hyper-nuclear matter, and ordinary nuclear matter. The modeled hybrid star confirms the mass, radius, and tidal deformability constraints of NSs.

### v3

@@
11institutetext: Institute for Theoretical Physics, University of Wrocław, Plac Maksa Borna 9, 50-204 Wrocław, Poland 22institutetext: Incubator of Scientific Excellence, Center for Simulations of Superdense Fluids, University of Wrocław, Plac Maksa Borna 9, 50-204 Wrocław, Poland 33institutetext: Frankfurt Institute for Advanced Studies, Giersch Science Center, D-60438 Frankfurt am Main, Germany 44institutetext: Alikhanyan National Science Laboratory (Yerevan Institute of Physics), Alikhanyan Brothers St 2, 0036 Yerevan, Armenia 55institutetext: Technische Universität Darmstadt, Fachbereich Physik, Institut für Kernphysik, Schlossgartenstraße 9, D-64289 Darmstadt, Germany 66institutetext: GSI Helmholtzzentrum für Schwerionenforschung GmbH, Theorie, Planckstraße 1, D-64291 Darmstadt, Germany
@@
Mahboubeh Shahrbaf email: mahboubeh.shahrbafmotlagh@uwr.edu.pl    Davood Rafiei Karkevandi email: Davood.Rafiei64@gmail.com    Alexander Ayriyan email: alexander.ayriyan@uwr.edu.pl    Stefan Typel
@@
Context. The presence of dark matter in neutron stars is of growing interest due to its potential impact on the structure and observable properties of these objects. Among the various candidates, the hypothetical sexaquark has emerged as a promising bosonic dark matter particle, potentially forming under extreme conditions in neutron star cores.
Aims. We investigate whether a hybrid neutron star model that includes hyperons, bosonic dark matter (in the form of sexaquarks), and deconfined quark matter can satisfy all current observational constraints. We particularly focus on identifying the range of sexaquark masses consistent with mass-radius measurements and the tidal deformability limit.
Methods. We used the DD2Y-T model for the hadronic phase, which includes hyperons, and a nonlocal Nambu–Jona-Lasinio model for the deconfined quark phase. The phase transition was modeled as a smooth crossover using the replacement interpolation construction method. Sexaquark-baryon interactions were introduced via an effective mass shift representing repulsion. We incorporated the full set of current observational data, including NICER measurements of PSRs J0437–4715 and newly published J0614–3329 data, and performed a Bayesian analysis to constrain the sexaquark mass.
Results. Our results show that the presence of the sexaquark softens the equation of state, enabling the hybrid model to satisfy both the radius and tidal deformability constraints around the canonical $1.4\,M_{\odot}$ neutron stars. We find that hybrid EOSs with a sexaquark mass around $1900$ MeV are in agreement with all available constraints, including those from HESS J1731–347 and PSR J0952–0607, which represent the lowest and highest mass neutron stars observed to date. The Bayesian analysis favors a sexaquark mass range of 1885188519351935 MeV, supporting the potential relevance of this exotic particle in neutron star interiors.
@@
Hyperons – Dark matter – Sexaquark – Hybrid Stars – Equation of State – Bayesian Analysis
@@ 1 Introduction
The accepted cosmological models that are presented and founded in general gelativity (GR) suggest that almost 27% of the matter-energy composition of the Universe exists in the form of enigmatic dark matter (DM; Bertone and Hooper, 2018; de Martino et al., 2020). Multi-messenger astrophysics is actively engaged in the search for and examination of new ideas concerning DM particles. These candidates are believed to be non-relativistic particles that exist beyond the Standard Model of particle physics. They interact weakly with ordinary baryonic matter (BM) solely through gravitational forces (Salucci and others, 2021; Arbey and Mahmoudi, 2021). In this context, weakly interacting massive particles are one of the notable DM candidates (Bertone and Tait, 2018; Arcadi et al., 2018). Moreover, hypothetical bosonic particles such as the axion (Marsh, 2016; Chadha-Day et al., 2022) and the sexaquark (S; (Farrar, 2022; Doser et al., 2023; Moore and Slatyer, 2024)) also exhibit the primary characteristics required to be considered potential DM candidates. Their interaction with the BM must be very weak, which is why they have remained undetected thus far. The mass of candidate DM particles spans a wide range, from approximately $10^{-22}$ eV to $10^{10}$ eV, and is based on the particular type of candidate (Cirelli et al., 2024). Moreover, their self-interaction is another key parameter of the DM model. Indeed, self-interacting DM has the potential to address several issues in cosmology and astrophysics, including ”core-cusp,” ”missing satellite,” and ”too big to fail” problems (Spergel and Steinhardt, 2000; Peter et al., 2013; Kaplinghat et al., 2016). Given these model parameters, DM particles must be incorporated into an equation of state (EOS), which is a fundamental tool for characterizing matter.
@@
It is believed that gravitationally stable astrophysical objects composed of DM could form in the Universe (Liebling and Palenzuela, 2023; Baryakhtar and others, 2022; Visinelli, 2021). In this context, they can exist as an object made entirely of DM, so-called dark stars (Narain et al., 2006; Maselli et al., 2017). Such objects include axion stars, boson stars (Eby et al., 2016; Visinelli et al., 2018; Pitz and Schaffner-Bielich, 2023), and stars that have a fraction of DM mixed with BM (Bramante and Raj, 2024). For the last type, compact objects such as black holes (Shakeri and Hajkarim, 2023), neutron stars (NSs; (Karkevandi et al., 2022; Shakeri and Karkevandi, 2024)), white dwarfs (Ryan and Radice, 2022), and even hypothetical quark (strange) stars (Ferreira and Fraga, 2023; Yang and Pi, 2024; Sedaghat et al., 2025; Liu et al., 2025a) offer a significant opportunity to probe DM indirectly. These objects are considered promising astrophysical environments where DM may reside, either within them or in a surrounding halo.
@@
Neutron stars, due to their extreme gravitational potential and high baryon density, have the potential to accumulate DM particles and are widely regarded as astrophysical laboratories for studying DM (Grippa et al., 2025). In recent years, motivated by new and significant observational data on NS properties, numerous studies have investigated the possible presence of DM within these objects (Ivanytskyi et al., 2020; Rafiei Karkevandi et al., 2021; Dengler et al., 2022; Rutherford et al., 2023; Karkevandi et al., 2024; Thakur et al., 2024a; Konstantinou, 2024; Routaray et al., 2025; Biesdorf et al., 2025; Arvikar et al., 2025). The influence of DM on observable features of NSs has been explored in order to place constraints on the parameter space of DM models (Sedrakian, 2016, 2019; Shakeri and Karkevandi, 2024; Giangrandi et al., 2023; Diedrichs et al., 2023; Ávila et al., 2024; Issifu et al., 2025; Cipriani et al., 2025; Scordino and Bombaci, 2025; Mukherjee et al., 2025). Furthermore, unusual measurements of NS mass and radius have prompted attempts to explain them in terms of mixed configurations, hereafter referred to as DM admixed NSs. In this direction, HESS J1731-347 (Sagun et al., 2023; Marzola et al., 2025) and XTE J1814-338 ultracompact stars (Pitz and Schaffner-Bielich, 2025; Yang et al., 2025; Lopes and Issifu, 2025) as well as very massive compact objects, such as the secondary ones detected in GW190814 and GW230529 events, are noteworthy cases (Barbat et al., 2024; Vikiaris et al., 2024).
@@
Concerning the formation of DM admixed NSs, several scenarios have been considered, each leading to various DM fractions within the star (Karkevandi et al., 2022). Among these, DM production inside NSs or during supernova explosions is a notable mechanism (Ellis et al., 2018; Nelson et al., 2019). An additional neutron decay channel into DM, potentially addressing the neutron decay anomaly, could result in the production of a significant fraction of fermionic DM inside NSs (Shirke et al., 2023, 2024; Thakur et al., 2024b; Das and Burgio, 2025). Furthermore, as a novel concept, the production of S, a deeply bound bosonic DM candidate, offers a promising mechanism for forming these mixed objects (Shahrbaf et al., 2022; Blaschke et al., 2022; Shahrbaf, 2023). This mechanism is the focus of this paper.
@@
Apart from the formation of these objects, modeling and investigation of DM admixed NSs are done based on the fact that DM and BM interact only through gravitational force or that there is a non-gravitational interaction between them. For the former, two-fluid Tolman-Oppenheimer-Volkoff equations describe the mixed object (Rafiei Karkevandi et al., 2021; Collier et al., 2022; Karkevandi et al., 2024; Rutherford et al., 2024; Shawqi and Morsink, 2024; Dengler et al., 2025; Liu et al., 2025b). Depending on the DM parameters, it can reside as a core inside an NS, or it may form an extended halo around the NS. For more detail on this first approach, we refer to the works done by Karkevandi et al. (2022); Shakeri and Karkevandi (2024) and the references therein. However, when a non-gravitational interaction is considered between BM and DM, single fluid Tolman-Oppenheimer-Volkoff equations are employed to describe the hydrostatic equilibrium of the object (Lenzi et al., 2023; Shahrbaf, 2023; Guha and Sen, 2024; Hajkarim et al., 2024; Kumar and Sotani, 2025; Klangburam and Pongkitivanichkul, 2025; Das et al., 2025). Therefore, DM production inside NSs, such as S, is investigated using the single-fluid approach, due to the non-gravitational interaction between DM particles and the baryonic medium (Shahrbaf et al., 2022). In this paper, we demonstrate how the presence of S affects the EOS, leading to changes in the observable properties of DM admixed NSs.
@@
It is worth mentioning that the EOS for strongly interacting matter in NSs is tightly bound by observational constraints. Therefore, efficient boundaries on the features of DM candidates have been established using recent advancements in observational astronomy, mainly facilitated by gravitational wave (GW) interferometers (Abbott and others, 2018; Dietrich et al., 2020; Pang and others, 2023; Ecker et al., 2025) and the Neutron Star Interior Composition ExploreR (NICER; (Miller and others, 2019, 2021; Riley and others, 2019, 2021; Vinciguerra and others, 2024; Salmi and others, 2024b, a; Dittmann et al., 2024; Choudhury and others, 2024; Mauviard and others, 2025)). In this study, mass, radius, and tidal deformability serve as the fundamental constraints used to investigate these DM admixed objects. Thus, we focus on high-precision mass–radius (M-R) measurements obtained through X-ray pulse profile modeling by the NICER telescope. We include the recent result for PSR J0437-4715, which reports a mass of $M=1.418\pm 0.037M_{\odot}$ and a radius of $R=11.36^{+0.95}_{-0.63}$ km Choudhury and others (2024). Importantly, we also incorporate the latest NICER measurement of PSR J0614-3329, which infers a radius of $R=10.29^{+1.01}_{-0.86}$ km for a mass of $M=1.44^{+0.06}_{-0.07}M_{\odot}$ (Mauviard and others, 2025). These new NICER results are important, as they support a softer EOS and add valuable boundary conditions for constraining models of dense matter in NSs.
@@
Moreover, the detection of GW signals arising from binary NS mergers, along with the subsequent determination of the tidal deformability parameter, have opened a new window for investigating the exotic internal structures of NSs (Abbott and others, 2017, 2020). Therefore, the tidal deformability constraint from the GW170817 event (Abbott and others, 2018), $\Lambda_{1.4}=190^{+390}_{-120}$, is also considered to obtain the allowed parameter space of the S bosonic DM model.
@@
The concept of a deeply bound S, an alternative term for the hexaquark family of hypothetical particles, was first introduced as a bosonic candidate for DM in references Farrar (2003, 2018, 2022); Doser et al. (2023). The quark composition of S is identical to that of the H-dibaryon (Jaffe, 1977), with both consisting of uuddss quarks. However, S exhibits significantly stronger binding compared to the H-dibaryon. If S represents a molecular state such as $\Lambda\Lambda$, the binding of two $\Lambda$ particles can exclusively occur through the exchange of color-neutral entities, such as mesons. However, when S comprises a complex system consisting of three colored diquarks, these entities engage in interactions primarily governed by the color force. This force is notably stronger than the meson exchange force, particularly when considering short-range distances.
It is important to note that the $S$ particle is a member of the wider family of hexaquark states. The first nontrivial hexaquark to be experimentally established was the $d^{*}(2380)$ dibaryon, discovered by the WASA-at-COSY Collaboration (Adlarson and others, 2011). This state has been investigated as an exotic degree of freedom in NSs. It has been shown that the $d^{*}(2380)$ could appear at baryon densities three to five times the saturation density, leading to a reduction in the maximum NS mass (Vidaña et al., 2018; Celi et al., 2024). More recent work Celi et al. (2025) has demonstrated that the inclusion of the $d^{*}(2380)$ softens the hadronic EOS and consequently delays the onset of deconfinement in a first-order phase transition. These studies highlight that an appropriate reparametrization of the hadronic models is required to remain consistent with astrophysical constraints when exotic dibaryons are included.
@@
The proposed S is a boson at a spin-color-flavor singlet and even parity state, and it is characterized by a zero electric charge, a baryon number of two, and a strangeness value of -2. The potential of the S particle to serve as DM depends on several parameters, such as its mass and its stability against dissociation into two baryons, which requires a small dissociation amplitude. Other important factors include its elastic scattering cross section with nucleons and its annihilation with antibaryons (or antisexaquarks) into baryons or other Standard Model final states. S may manifest itself as a deeply bound state, displaying a mass sufficiently low to either be absolutely or effectively stable against decay. Our main focus in this study is the mass of the S. Notably, all measurements show that its mass is less than 2230 MeV, the threshold for producing two $\Lambda$ baryons. This suggests that the S may be a bound state (Jaffe, 1977; Kodama et al., 1994; Azizi et al., 2020; Gross et al., 2018; Shahrbaf et al., 2022; Evans and Ward, 2023).
@@
However, Lattice QCD calculations suggest that H-dibaryon is an unbound particle (Inoue et al., 2011; Beane and others, 2011; Green et al., 2021; Shanahan et al., 2011), but Farrar (2022) assumes S as a deeply bound state that is difficult to detect in lattice calculations, making it different from the loosely bound H-dibaryon. For S to be a viable DM candidate, both its possible decay channels and any Standard Model particle decays involving the S must either be kinematically forbidden or highly suppressed. If the S particle were to decay, the energetically most favorable channel would produce two protons and two electrons, with a total mass of approximately $1878$ MeV. Consequently, the S particle would be absolutely stable if its mass lies below this threshold, i.e., $m_{\text{S}}<1878$ MeV. From an experimental point of view, the existence of S with a mass between $2054$ MeV and $2223.7$ MeV has been ruled out (Takahashi and others, 2001). Additionally, Farrar and Wang (2023) argues that if the S mass equals the combined mass of a lambda hyperon, a proton, and an electron ($m_{\text{S}}=2054$ MeV), its decay into two nucleons would involve a doubly weak interaction, resulting in a lifetime longer than the age of the Universe. The constraints on the mass of the S have been extensively discussed in Farrar and Wang (2023) and the references therein. Meanwhile, Gal (2024) have demonstrated that the observation of double-$\Lambda$ hypernuclei decaying into single-$\Lambda$ hypernuclei does not exclude the existence of a deeply bound S within the mass range of $2m_{n}$ to $m_{\Lambda}+m_{n}$. However, its $\Delta S=2$ decay lifetime ($\tau(H\rightarrow nn)$) is estimated to be shorter than the timescale required for it to be a viable DM candidate. Moreover, the ratio of densities between S DM and ordinary BM in the Universe, which can be estimated through considerations of statistical equilibrium within the quark-gluon plasma, aligns with observational data (Farrar, 2022). Under the above-mentioned conditions, we assume that the S has the potential to be a DM candidate and examine its mass based on observational constraints from NSs. A mechanism for producing such an uuddss state at rest through the formation of antiprotonic atoms has been proposed in Doser et al. (2023). Ultimately, lattice QCD can be used to study the interactions and decay channels of the S, offering deeper insight into its properties.
@@
Complementary to the S DM hypothesis, heavy-ion collision experiments have demonstrated that hyperons can be abundantly produced in high-density environments. These findings are significant for understanding the behavior of dense nuclear matter, as they provide empirical evidence supporting the inclusion of hyperons in the EOS of matter under extreme conditions (Abelev and others, 2010; Rappold and others, 2015; Acharya and others, 2022; Leung, 2022; Acharya and others, 2023). Therefore, including hyperons in NS EOS models is essential for a comprehensive understanding of their structure, stability, and evolution, aligning theoretical predictions with observational data. Moreover, given that our DM candidate carries double strangeness, it is natural to include hyperons in the hadronic phase as well. In this study, the S production is established using a generalized relativistic density functional approach known as DD2Y-T (Stone et al., 2021). This hadronic model contains all $J=1/2$ octet baryons, namely, the neutron, proton, $\Lambda$, $\Sigma$, and $\Xi$ hyperons, giving a more reliable description of the high-density matter inside NSs. The EOS for the entire NS, including both the core and the crust, is constructed consistently using the DD2 crust model as in Typel and Shlomo (2024), which ensures a smooth and thermodynamically consistent matching between the outer layers and the dense core. In addition to the strongly interacting baryonic components, the EOS must also account for leptonic degrees of freedom - specifically, electrons and muons - due to the requirements of charge neutrality and beta equilibrium in NS matter. These leptons are treated as non-interacting, relativistic Fermi gases and play a crucial role in balancing the electric charge introduced by the presence of protons and charged hyperons.
@@
Furthermore, from a phenomenological point of view, a peak in the square of the speed of sound in QCD matter has been observed, which is associated with the percolation threshold and the phase transition to deconfined QM (Satz, 1998; Magas and Satz, 2003; Castorina et al., 2009; Braun-Munzinger et al., 2015; Fukushima et al., 2020; Marczenko et al., 2023). These findings support the idea that the core of NSs may consist not only of hadronic matter but also deconfined QM. As a result, it is highly possible to encounter a phase transition to QM in the core of NS (Li et al., 2023; Gholami et al., 2025; Li et al., 2025; Lindblom et al., 2025). Therefore, our model is connected to an advanced QM model named the nonlocal Nambu-Jona-Lasinio (nlNJL) model, which is based on the crossover phase transition through a novel method.
Consequently, we modeled hybrid stars with a dense core of deconfined QM surrounded by layers of bosonic S DM, hypernuclear matter, and ordinary nuclear matter. Our aim is to examine whether a unified EOS incorporating these exotic components can satisfy current multi-messenger constraints on NSs while also placing constraints on the upper bound of the bosonic DM mass from astrophysical observations.